OpenStudy (prettygirlnikki):

does anyone know how to multiply matrices ? I've been trying for the longest and i still don't understand it. If you know, is it possible for you to explain it to me so that i can understand it better?

OpenStudy (anonymous):

Not sure. I just kinda knew it. I think of it as equations with the lines in order all ready to multiply. Sorry, if that doesn't help (which I wouldn't expect it to xD) lol

OpenStudy (shadowlegendx):

I use a calculator like the lazy nub I am

OpenStudy (shadowlegendx):

Iʻm sure Kainui will show you how to do it by hand tho.

OpenStudy (kainui):

@prettygirlnikki in matrix notation, the dot product of two vectors looks like this: \[ \begin{bmatrix} a & b \end{bmatrix} \begin{bmatrix} x \\ y \end{bmatrix} = ax+by\] When you multiply two matrices, you're taking the dot product of the every row vector in the left matrix with every column vector in the right matrix.

OpenStudy (prettygirlnikki):

okay, im going to try it. thanks

OpenStudy (greatlife44):

@Kainui killing it once more

OpenStudy (greatlife44):

great work as always